# Fast Fourier Transform periodic interpolation method for superposition sums in a periodic unit cell

### v1

@@ 3.3 Evaluation of the near-zone PSP
To evaluate the near-zone PSP component, $u_{near}$, we note that for $i_{d}=0$, we only consider the zeroth order unit cell, which is equivalent to a non-periodic problem. The evaluation of $u_{near}$ for a free-space standalone problem can be efficiently handled using various fast computational approaches, such as the fast multipole like methods [5, 6] or interpolation-based approaches [8, 9, 10]. For $i_{d}>0$, neighbor images in the near-zone evaluation are included and they require modifying the fast evaluation algorithms. Here, we use the box-adaptive integral method of [11, 25], which is also related to the pre-corrected FFT method [14] and adaptive integral method [12, 13]. We modify this approach to efficiently account for the inclusion of the periodic images. The idea is similar to the far-zone procedure in Sec. 3.2 in its first three steps, but it also adds another modified correction step.
@@
Step 4: Error correction: The ability to perform the projection/interpolation in steps 1 and 3 rely on the assumption of slow variations of the Green’s function. For the far-zone evaluation, $G^{p}_{far}$ indeed varies slowly in the domain of the zeroth unit cell because it excludes the direct and near-zone image interactions between the sources and observers. The near-zone component with $G^{p}_{near}$, however, involves direct and near-image interactions, which result in rapid spatial variations, e.g., the rate of such spatial variations is unbounded when the source and observer coincide. Therefore, the interpolations for the nearby source-observer pair contributions are highly inaccurate, and they need to be corrected. To eliminate these errors, we define an error correction range $D^{ER}$ for each box, determined by the uniform grid. This range is of the same order as $max\{\Delta_{x},\Delta_{y},\Delta_{z}\}$. For all observer points in a box, the corrections are performed in the error-correction region, $\Omega^{ER}(\mathbf{r}_{m})$, for all the non-uniform sources in the same box and a certain number of surrounding boxes that are withing the $D^{ER}$ distance from the box. The correction involves subtracting the contributions due to steps 1-3 and adding the exact point-to-point superposition sums via appropriate contributions in Eq. (1). The error correction procedure can be expressed as
@@
where the error is corrected in the region $\Omega^{ER}(\mathbf{r}_{m})$, i.e., for any point $\mathbf{r}$ within this region, $|\mathbf{r}-\mathbf{r}_{m}|\leq D^{ER}$. Similar to Eq. (11) and Eq. (13), the $\omega^{s}_{nufft}$ and $\omega^{o}_{nufft}$ correspond to the interpolation matrices containing the weights for the above step 1 and 2. The function $G_{grid}$ represents the near-zone Green’s function between the source and observer points obtained via the grid interactions with interpolation, which needs to be subtracted in the error correction process. While the procedure in Eq. (3.3) can be efficient, it still can significantly increase the computational cost due to the need to account for all the $i_{d}$ near-zone images via $G^{p}_{near}$. For example, in the 3D periodicity case, $i_{d}=125$ when including 2 surrounding boxes per each dimension, which leads to a high additional computational cost.
@@
This computational cost can be much reduced by including only a single extra periodic image per periodicity dimension. This is possible due to the fact that $\{\Delta_{x},\Delta_{y},\Delta_{z}\}$ are typically much smaller than $\{D_{x},D_{y},D_{z}\}$, so that most of the images correspond to a large spatial separation. For example, in Fig. 4 showing an example for a 1D $x$-dimension periodicity case, for the sources in the internal boxes of the zeroth (middle) unit cell, only the $G_{0}(\mathbf{r}_{m}-\mathbf{r}_{n})$ component of the $G^{p}_{near}(\mathbf{r}_{m}-\mathbf{r}_{n})$ can be used for the near-zone correction. On the other hand, for the left box, which is next to the left edge of the zeroth unit cell (the left green box with red dot and arrows), in addition to the $G_{0}(\mathbf{r}_{m}-\mathbf{r}_{n})$ component, the component of $G_{0}(\mathbf{r}_{m}-\mathbf{r}_{n}-L_{x}\hat{x})$ of $G^{p}_{near}(\mathbf{r}_{m}-\mathbf{r}_{n})$ needs to be accounted for. The latter component corresponds to the sources from the right box (right green box with a black dot and arrows) contributing to the left box observers via the $i_{x}=1$ image. These sources are equivalent to the sources in the right box of the left image cell (the dashed green box with a grey dot) contributing to the left box observers via the image Green’s function $G_{0}(\mathbf{r}_{m}-\mathbf{r}_{n})$, i.e., they are equivalent to the geometrically close boxes. This procedure is generalized to 2D and 3D periodicities by including corresponding required images. The computational cost of this procedure is only slightly higher than that for a non-periodic case, in which only the $G_{0}(\mathbf{r}_{m}-\mathbf{r}_{n})$ contribution is required. The cost increase is associated with the need to include additional images at the sources associated with the boxes located at the corners, edges, and sides of the computational domain. There is a small number of such boxes as compared to the total number of near-zone images, so that the computational cost increase is insignificant. The overall computational cost of this step is of $O(N)$.
@@ 4 Numerical results
We start by showing results for the calculated potential due to a source distribution (Fig. 5). The structure is 1D periodic in the $x$-direction with the periodicity of $L_{x}=1$ and the unit cell comprises a section of coaxial cable along the $x$-axis that extends through the entire period. The inner and outer radius is $r_{1}=1$ and $r_{2}=2$, respectively. Negative and positive sources (e.g., charges) are distributed uniformly on the inner surface with charge density $\rho_{1}=-1$ and $\rho_{2}=1/2$, respectively, resulting in a neutral unit cell. We first calculate the potential on the central $x$ axis for a non-periodic case, as shown in Fig. 5 and we observe significant edge effects in terms of the potential spatial dependence. We then use the FFT-PIM with a 1D periodicity, making it into an infinitely long coaxial source distribution. With the static condition and no phase shift, we find that the potential becomes near zero everywhere (blue curve in Fig. 5), the edge effect is eliminated, which agrees with the analytical result. We also calculate the potential for the dynamic case with the wavenumber $k_{0}=1$ and a complex phase shift $k_{x0}=1-j$, in which case the potential becomes complex with non-zero magnitude and corresponding real/imaginary parts (black, red, and green curves in Fig. 5).
Figure 5: Potential on the $x$ axis of a coaxial structure . (a) Illustration of coaxial structure unit cell ($L=1$) and potential of the standalone unit cell. Inner radius $r_{1}=1$ with negative line charge density $\rho_{1}=-1$, outer radius $r_{2}=2$ with positive line charge density $\rho_{2}=1/2$. (b) PSP with a 1D periodicity along the $x$ axis with $L_{x}$ = 1. The blue line is the magnitude of PSP for the static case with $k_{0}=k_{x0}=0$. The red, green and black lines are real part, imaginary part, and magnitude of PSP for the dynamic case with $k_{0}=1$, $k_{x0}=1-j$.
@@ 4.1 Error analysis
To test the error of the FFT-PIM, we use an infinite long 1D bar along the $x$ axis, and the target object is a cube of the size $D_{x},D_{y},D_{z}=50$ with 1D $x$-direction periodicity with $L_{x}=D_{x}=50$. We assume a static case without a phase shift, i.e., $k=k_{x0}=0$. We mesh the domain with a tetrahedral mesh and the total number of the mesh points (vertices) is $N=7189$. In order to calculate the error, we randomly set 7000 points with finite values as source points while the rest of the sources are set to 00, and make sure that $\sum_{n=1}^{N}q_{n}=0$. The rest of the points are considered as observer points, and the location of the observer points are chosen such that they are spatially shifted as compared to the source points such that the PGF in Eq. (5a) converges for all source-observer pairs when the direct sum is used. We compare the results for errors between the FFT-PIM and the direct computation via Eq. (5a), which is equivalent to Eq. (3a) as shown in Sec. 2. We tested the errors at the observer points with different interpolation orders, uniform grid size, and number of subtracted near-zone images. The results are shown in Fig. 6. We find that with cubic interpolation order and a uniform grid of size $N_{gx},N_{gy},N_{gz}=10$, by subtracting the zeroth order and only the nearest neighbors ($i_{d}=1$), we can achieve an error at the level of $10^{-3}$, which is sufficient for many practical cases. The error is further reduced by increasing the number of grid points, order, and the near-zone images.
Figure 6: Error of the far-zone PSP for a 1D periodicity along the $x$-axis. (a) Error for 1st, 3rd, and 6th order interpolation with varying uniform grid sizes with $i_{d}=1$. (b) Error of cubic interpolation with 1000 sparse grid points and different numbers of subtracted near-zone unit cells.
@@ 4.2 Computational performance analysis
Figure 7: Preprocessing time for the no-periodicity case, 1D periodicity case, and for the difference between the two, i.e., for the overhead due to the periodicity.
Figure 8: Execution time of evaluating the potential for the standalone case and for PSP. (a) Execution time on a single-core CPU with (red line) and without (blue line) periodicity, highlighting a marginal difference (black dashed line) between the two configurations. (b) Execution time on GPU with (red line) and without (blue line) periodicity, where the difference (black dashed line) is negligible.
To benchmark the performance of the FFT-PIM, we conducted a set of tests on a workstation equipped with AMD 5950X and NVIDIA 3080 Ti, with the CPU tests utilizing a single core. The results are shown for a 3D periodicity static problem without phase shifts. The target object is a cube with dimensions $D_{x},D_{y},D_{z}=100$ and $L_{x,y,z}=D_{x,y,z}+1$. This problem has $L_{x,y,z}\approx D_{x,y,z}$ and it represents the maximal computational complexity as compared to 1D and 2D periodicity cases. The configuration of the far-zone calculation is set as cubic interpolation with uniform grid size of $N_{gx},N_{gy},N_{gz}=10$, and $i_{d}=1$. The near-zone is handled using the procedures from the FastMag micromagnetic simulator [22, 25], which is modified to allow for an efficient handling of periodicities, as outlined in Sec. 3.3.
@@
Fig. 8 shows the execution time for a single evaluation of PSP on CPU and GPU. When compared to a non-periodic problem, the major execution time increase is due to the evaluation of the far-zone PSP component, which can be performed separately from the near-zone evaluation. The presence or absence of periodicity has minimal impact on the near-zone evaluation since all Green’s function values are tabulated during the pre-processing step. Consequently, in the single-core CPU benchmark, only a minor difference is observed between configurations with and without periodicity. In contrast, in the GPU benchmark, NVIDIA CUDA multi-stream concurrency enables simultaneous implementation of both near-zone and far-zone evaluations. This approach effectively eliminates the overhead induced by the far-zone evaluation, which is substantially slower, resulting in negligible observable differences between cases with and without periodicity, even in the 3D periodicity cases. The 1D and 2D periodicity cases exhibit a similar performance, indicating a highly efficient implementation for all 1D, 2D, 3D periodicity types.
@@ 5 Summary and discussion
We introduced an efficient and flexible FFT-PIM for computing PSP. FFT-PIM can be used for a wide variety of problem types with the same numerical implementation. It allows computing PSP for a non-uniform source distribution, works for arbitrary 1D, 2D, and 3D periodicities, can operate with or without a phase shift between the periodic boundaries, and is applicable to dynamic and static problems. A requirement is imposed for using FFT-PIM for static problems with a vanishing phase shift, which is that the source must be neutral within the unit cell. Such a requirement is natural to many practical problems, such as problems dealing with electric dipoles, magnetization, and molecular structures.
@@
FFT-PIM is based on a superposition sum between the source distribution and PGF. PGF, which originally is defined as an infinite sum that does not have a rapid convergence, can be represented as an exponentially convergent sum in terms of spectral series expansions, which are given for the dynamic case and static case with a periodic phase shift as well as for a static case with no phase shift. The expansions converge under the assumption of a sufficient separation between the source and observer points. To lead to a fast superposition sum computation, PGF and the corresponding PSP are represented in terms of the near- and far-zone components. The near-zone component includes a finite, typically small, number of periodic images, whereas the far-zone component includes all the rest of an infinite number of images. The far-zone component is recognized as a slowly varying function of the spatial coordinates, which allows computing the far-zone PSP component by first computing it on a sparse uniform grid, i.e., at a small number of points, and then interpolating it to all the required non-uniform observation points by local interpolation. This process involves defining shifted source and observer grids to allow for a rapidly convergence of the spectral sums for PGF. This process requires a small number of operations and low computational cost, while allowing for a rapid converge. The near-zone PSP component can be evaluated by any fast method, but it required accounting for the additional near-zone images, which may substantially increase the computational cost and memory consumption. An approach based on the box-adaptive adaptive integral method is presented that allows evaluating the superposition sum for the far-zone component rapidly based on including essential images only at the required locations near the periodic boundaries. The result is an approach that has an overall computational cost of $O(N\log N)$ and memory consumption of $O(N)$. The presented results demonstrate the high convergence, accuracy, and computational performance of the FFT-PIM.

### v2

@@ 3.3 Evaluation of the near-zone PSP
To evaluate the near-zone PSP component, $u_{near}$, we note that for $i_{d}=0$, we only consider the zeroth order unit cell, which is equivalent to a non-periodic problem. The evaluation of $u_{near}$ for a free-space non-periodic problem can be efficiently handled using various fast computational approaches, such as the fast multipole like methods [5, 6] or interpolation-based approaches [8, 9, 10]. For $i_{d}>0$, neighbor images in the near-zone evaluation are included and they require modifying the fast evaluation algorithms. Here, we use the box-adaptive integral method of [11, 36], which is also related to the pre-corrected FFT method [14] and adaptive integral method [12, 13]. We modify this approach to efficiently account for the inclusion of the periodic images. The idea is similar to the far-zone procedure in Sec. 3.2 in its first three steps, but it also adds another modified correction step.
@@
Step 4: Error correction: The ability to perform the projection/interpolation in steps 1 and 3 rely on the assumption of slow variations of the Green’s function. For the far-zone evaluation, $G^{p}_{far}$ indeed varies slowly in the domain of the zeroth unit cell because it excludes the direct and near-zone image interactions between the sources and observers. The near-zone component with $G^{p}_{near}$, however, involves direct and near-image interactions, which result in rapid spatial variations, e.g., the rate of such spatial variations is unbounded when the source and observer coincide. Therefore, the interpolations for the nearby source-observer pair contributions are highly inaccurate, and they need to be corrected. To eliminate these errors, we define an error correction range $D^{ER}$ for each box (defined by $\Delta_{x}$, $\Delta_{y}$ and $\Delta_{z}$ in Eq. (8c)), determined by the uniform grid. This range is of the same order as $max\{\Delta_{x},\Delta_{y},\Delta_{z}\}$. For all observer points in a box, the corrections are performed in the error-correction region, $\Omega^{ER}(\mathbf{r}_{m})$, for all the non-uniform sources in the same box and a certain number of surrounding boxes that are withing the $D^{ER}$ distance from the box. The correction involves subtracting the contributions due to steps 1-3 and adding the exact point-to-point superposition sums via appropriate contributions in Eq. (1). The error correction procedure can be expressed as
@@
where the error is corrected in the region $\Omega^{ER}(\mathbf{r}_{m})$, i.e., for any point $\mathbf{r}$ within this region, $|\mathbf{r}-\mathbf{r}_{m}|\leq D^{ER}$. Similar to Eq. (11) and Eq. (13), the $\omega^{s}_{nufft}$ and $\omega^{o}_{nufft}$ correspond to the interpolation matrices containing the weights for the above step 1 and 2. The function $G_{grid}$ represents the near-zone Green’s function between the source and observer points obtained via the grid interactions with interpolation, which needs to be subtracted in the error correction process. While the procedure in Eq. (14) can be efficient, it still can significantly increase the computational cost due to the need to account for all the near-zone images via $G^{p}_{near}$. For example, in the 3D periodicity case, when including 2 surrounding boxes per each dimension ($i_{d}=2$), there are 125 ($(2\times i_{d}+1)^{3}$) images in total, which leads to a high additional computational cost.
@@
This computational cost can be much reduced by including only a single extra periodic image per periodicity dimension. This is possible due to the fact that $\{\Delta_{x},\Delta_{y},\Delta_{z}\}$ are typically much smaller than $\{D_{x},D_{y},D_{z}\}$, so that most of the images correspond to a large spatial separation. For example, in Fig. 4 showing an example for a 1D $x$-dimension periodicity case, for the sources in the internal boxes of the zeroth (middle) unit cell, only the $G_{0}(\mathbf{r}_{m}-\mathbf{r}_{n})$ component of the $G^{p}_{near}(\mathbf{r}_{m}-\mathbf{r}_{n})$ can be used for the near-zone correction. On the other hand, for the left box, which is next to the left edge of the zeroth unit cell (the left green box with red dot and arrows), in addition to the $G_{0}(\mathbf{r}_{m}-\mathbf{r}_{n})$ component, the component of $G_{0}(\mathbf{r}_{m}-\mathbf{r}_{n}-L_{x}\mathbf{\hat{x}})$ of $G^{p}_{near}(\mathbf{r}_{m}-\mathbf{r}_{n})$ needs to be accounted for. The latter component corresponds to the sources from the right box (right green box with a black dot and arrows) contributing to the left box observers via the $i_{x}=1$ image. These sources are equivalent to the sources in the right box of the left image cell (the dashed green box with a grey dot) contributing to the left box observers via the image Green’s function $G_{0}(\mathbf{r}_{m}-\mathbf{r}_{n})$, i.e., they are equivalent to the geometrically close boxes. This procedure is generalized to 2D and 3D periodicities by including corresponding required images. The computational cost of this procedure is only slightly higher than that for a non-periodic case, in which only the $G_{0}(\mathbf{r}_{m}-\mathbf{r}_{n})$ contribution is required. The cost increase is associated with the need to include additional images at the sources associated with the boxes located at the corners, edges, and sides of the computational domain. There is a small number of such boxes as compared to the total number of near-zone images, so that the computational cost increase is insignificant. The overall computational cost of this step is of $O(N)$.
@@ 4 Numerical results
We start by showing results for the calculated potential due to a source distribution (Fig. 5). The structure is 1D periodic in the $x$-direction with the periodicity of $L_{x}=1$ and the unit cell comprises a section of coaxial cable along the $x$-axis that extends through the entire period. The inner and outer radius is $r_{1}=1$ and $r_{2}=2$, respectively. Negative and positive sources (e.g., charges) are distributed uniformly on the inner surface with charge density $\rho_{1}=-1$ and $\rho_{2}=1/2$, respectively, resulting in a neutral unit cell. We first calculate the potential on the central $x$ axis for a non-periodic case, as shown in Fig. 5 and we observe significant edge effects in terms of the potential spatial dependence. We then use the FFT-PIM with a 1D periodicity, making it into an infinitely long coaxial source distribution. For the NPSP case, we find that the potential becomes near zero everywhere (blue curve in Fig. 5), the edge effect is eliminated, which agrees with the analytical result. We also calculate the potential for the dynamic case with the wavenumber $k_{0}=1$ and a complex phase shift $k_{x0}=1-j$, in which case the potential becomes complex with non-zero magnitude and corresponding real/imaginary parts (black, red, and green curves in Fig. 5).
Figure 5: Potential on the $x$ axis of a coaxial structure . (a) Illustration of coaxial structure unit cell ($L=1$) and potential of the non-periodic unit cell. Inner radius $r_{1}=1$ with negative line charge density $\rho_{1}=-1$, outer radius $r_{2}=2$ with positive line charge density $\rho_{2}=1/2$. (b) PSP with a 1D periodicity along the $x$ axis with $L_{x}$ = 1. The blue line is the magnitude of PSP for the NPSP case with $k_{0}=k_{x0}=0$. The red, green and black lines are real part, imaginary part, and magnitude of PSP for the dynamic case with $k_{0}=1$, $k_{x0}=1-j$.
@@ 4.1 Error analysis
To test the error of the FFT-PIM, we use an infinite long 1D bar along the $x$ axis, and the target object is a cube of the size $D_{x},D_{y},D_{z}=50$ with 1D $x$-direction periodicity with $L_{x}=D_{x}=50$. We assume a NPSP case, i.e., $k=k_{x0}=0$. We mesh the cube with a tetrahedral mesh and the total number of the mesh points (vertices) is $N=7189$. In order to calculate the error, we randomly set 7000 mesh points with finite values as source points while the rest 189 points are set to 00, and make sure that $\sum_{l=1}^{N}q_{l}=0$. The 189 mesh points set to 00 are considered as observer points, and the locations of these observer points are chosen such that they are spatially shifted as compared to the source points such that the PGF in Eq. (5a) converges for all source-observer pairs when the direct superposition sum is used. We compare the results for relative errors between the FFT-PIM and the direct computation via Eq. (5a), which is equivalent to Eq. (3a) as shown in Sec. 2. We tested the relative errors at the observer points with different interpolation orders, uniform grid size, and number of subtracted near-zone images. The results are shown in Fig. 6. We find that with cubic interpolation order and a uniform grid of size $N_{gx},N_{gy},N_{gz}=10$, by subtracting the zeroth order and only the nearest neighbors ($i_{d}=1$), we can achieve an error at the level of $10^{-3}$, which is sufficient for many practical cases. The relative error is further reduced by increasing the number of grid points, order, and the near-zone images.
Figure 6: Relative error of the far-zone PSP for a NPSP case with a 1D periodicity along the $x$-axis. (a) Relative error for 1st, 3rd, and 6th order interpolation with varying uniform grid sizes with $i_{d}=1$. (b) Relative error of cubic interpolation with 1000 sparse grid points and different numbers of subtracted near-zone unit cells.
@@ 4.2 Computational performance analysis
Figure 7: Preprocessing time for the non-periodic case, NPSP case with 3D periodicity, and for their difference.
Figure 8: Execution time versus $N$ of evaluating the non-periodic potential and NPSP PSP with 3D periodicity on (a) eight-core CPU and (b) on GPU.
$N$ Near-zone grid size
3096 $17^{3}$
12175 $25^{3}$
53601 $41^{3}$
92233 $49^{3}$
177973 $59^{3}$
418308 $77^{3}$
1391742 $111^{3}$
Table 1: Grid size for near-zone PSP component evaluation. The near-zone grid is much larger than the far-zone grid when the problem size $N$ is large. Second-order projection/interpolation for near-zone evaluation is used.
Figure 9: Execution time versus the relative error for the near-zone PSP component on CPU and GPU for a 3D NPSP case with $N=53K$ and $N=510K$.
To benchmark the performance of the FFT-PIM, we conducted a set of tests on Purdue Anvil cluster at Rosen Center for Advanced Computing (RCAC) in Purdue University. CPU tests were run on eight cores of AMD Epyc ”Milan” processors and GPU tests were run on Nvidia A100. The results are shown for a 3D NPSP problem. The target object is a cube with dimensions $D_{x},D_{y},D_{z}=100$ and $L_{x,y,z}=D_{x,y,z}+1$. This problem has $L_{x,y,z}\approx D_{x,y,z}$ and it represents the maximal computational complexity as compared to 1D and 2D periodicity cases. The configuration of the far-zone calculation is set as cubic interpolation with uniform grid size of $N_{gx},N_{gy},N_{gz}=10$, and $i_{d}=1$. The near-zone is handled using the procedures from the FastMag micromagnetic simulator [27, 36] for non-periodic and periodic cases. When periodicity presents, it is modified to allow for an efficient handling of periodicities, as outlined in Sec. 3.3.
@@
Fig. 8 shows the execution time for a single evaluation of PSP on CPU and GPU. When compared to a non-periodic problem, the major execution time increase is due to the evaluation of the far-zone PSP component, which can be performed separately from the near-zone evaluation. The presence or absence of periodicity has minimal impact on the near-zone evaluation since all Green’s function values are tabulated during the pre-processing step. Consequently, in the eight-core CPU benchmark, only a minor difference ($<15\%$) is observed between configurations with and without periodicity for moderately large case where $N>10^{4}$. Furthermore, in the GPU benchmark, NVIDIA CUDA multi-stream concurrency enables simultaneous implementation of both near-zone and far-zone evaluations. This approach effectively eliminates the overhead induced by the far-zone evaluation, which is already marginal in eight-core CPU execution, resulting in a small differences ($<5\%$) between cases with and without periodicity, including the 3D periodicity cases. The execution performance on GPU is generally 20-30X faster than that on eight-core CPU for the tested sizes, which is equivalent to around $\sim$200X speedup between GPU and single CPU core. The 1D and 2D periodicity cases exhibit a similar performance, indicating a highly efficient implementation for all 1D, 2D, 3D periodicity types. To further present differences between the near-zone and far-zone component, we show the grid size of the near-zone PSP component in Table 1. For most practical problem sizes, the near-zone grid in Table 1 is much denser than the far-zone grid.
From Fig. 8 we observe that the near-zone component PSP dominates the computational load for the 1D, 2D, and 3D periodicity cases. Similarly, Figure 6 reveals that the error associated with the far-zone PSP component can be reduced with minimal adjustments to the parameters and limited overhead. This observation suggests that the near-zone PSP component is a critical factor in determining the balance between overall performance and error rates. To further explore this relationship, we conducted a series of benchmarks with a cube meshed in its volume, containing $N=53,601$ points for a coarser mesh and $N=510,643$ for a finer mesh. The results are shown in Fig. 9. The performance shown here demonstrate a good performance compared to available solutions [16].
Figure 10: (a)Preprocessing time and (b) execution time versus $N$ of evaluating the PSP for a 3D NPSP and high-frequency dynamic cases for the relative errors of $31e-$ 3 and $51e-$ 5.
Figure 10 compares the preprocessing and computational time of the NPSP and dynamic cases for a 3D periodicity of $L_{x}=101$, $L_{y}=101$, $L_{z}=101$ for relative errors of $10^{-3}$ and $10^{-5}$. The NPSP case parameters are as those in Fig. 8 and Table. 1 for error level $10^{-3}$, and we double the near-zone grid size for error level $10^{-5}$. The dynamic cases have phase shifts of $k_{x0}=0.01-0.01j$, $k_{y0}=0.01-0.01j$, $k_{z0}=0.01-0.01j$. For the dynamic case, the wavenumber $k_{0}$ was chosen such that the average distance between the sources/observers was $\lambda/20$. This choice is typical when considering high-frequency dynamic problems and it allows resolving the wave process spatial variations. For the largest considered $N$, this choice results in the domain size of $L_{x}=L_{y}=L_{z}=5\lambda$. We find that the computational time for the low- and high-frequency cases are close to each other and they scale similarly as that for the non-periodic static case. The computational time of the dynamic cases is around 2 times greater, which is related to performing complex-valued versus real-valued operations.
@@ 5 Summary and discussion
We introduced an efficient and flexible FFT-PIM for computing PSP. FFT-PIM can be used for a wide variety of problem types with the same numerical implementation. It allows computing PSP for a non-uniform source distribution, works for arbitrary 1D, 2D, and 3D periodicities, can operate with or without a phase shift between the periodic boundaries, and is applicable to dynamic and static problems. A requirement is imposed for using FFT-PIM for NPSP problems, which is that the source must be neutral within the unit cell. Such a requirement is natural to many practical problems, such as problems dealing with electric dipoles, magnetization, and molecular structures.
@@
FFT-PIM is based on a superposition sum between the source distribution and PGF. PGF, which originally is defined as an infinite sum that does not have a rapid convergence, can be represented as an exponentially convergent sum in terms of spectral series expansions, which are given for the dynamic case and static case with and without a periodic phase shift. The expansions converge under the assumption of a sufficient separation between the source and observer points. To lead to a fast superposition sum computation, PGF and the corresponding PSP are represented in terms of the near- and far-zone components. The near-zone component includes a finite, typically small, number of periodic images, whereas the far-zone component includes all the rest of an infinite number of images. The far-zone component is recognized as a slowly varying function of the spatial coordinates, which allows computing the far-zone PSP component by first computing it on a sparse uniform grid, i.e., at a small number of points, and then interpolating it to all the required non-uniform observation points by local interpolation. This process involves defining shifted source and observer grids to allow for a rapidly convergence of the spectral sums for PGF. This process requires a small number of operations and low computational cost, while allowing for a rapid converge. The near-zone PSP component can be evaluated by any fast method, but it required accounting for the additional near-zone images, which may substantially increase the computational cost and memory consumption. An approach based on the box-adaptive adaptive integral method is presented that allows evaluating the superposition sum for the far-zone component rapidly based on including essential images only at the required locations near the periodic boundaries. The result is an approach that has an overall computational cost of $O(N\log N)$ and memory consumption of $O(N)$. The presented results demonstrate the high convergence, accuracy, and computational performance of the FFT-PIM.
We note that FFT-PIM can be regarded as an extension of the fast periodic interpolation method [26] and the non-periodic box-adaptive integral method [36] or pre-corrected FFT method [14]. With respect to the fast periodic interpolation method [26], the extensions are in enabling both dynamic and static cases with any phase shifts or no phase shift as well as in enabling the complete fields computation, including fast near-zone computations via an FFT-based method without a need of extending the computational domain. With respect to the non-periodic box-adaptive integral method [36] or pre-corrected FFT method [14], FFT-PIM allows considering problems with 1D, 2D, and 3D periodicities with almost the same computational cost as the non-periodic approaches. We also note that the approach of separating the computations into the near- and far-zone components is related to other fast evaluation techniques [14, 15, 16, 17, 18, 26, 36] for periodic and non-periodic domains. Contributions of this work are in the efficient evaluation of the far-zone component via the properly chosen uniform grids and Green’s function evaluation as well as in the adaptation of the near-zone component computation via an FFT-based method.
While FFT-PIM is a powerful method, it also has some limitations. In particular, for high-frequency problems with electrically large periodicities, the computation of PGF may become slow. This time may be reduced by using alternative methods for computing PGF, e.g., see Refs. [32] and [37]. Additionally, using FFTs mean that the computations are done on the entire FFT grid, which may be inefficient for problems of curved linear or surface domains or in cases of highly non-uniform domains with dense constellations of sources and observers in certain locations. For such problems using methods such as Fast Multiple Method [5], [38] or Non-Uniform Grid Interpolation method [9] can be efficient. These methods can substitute the FFT-based approach for the near-zone component and the ideas presented here can be adapted to extend these methods to account for the additional images required due to periodicity.
FFT-PIM can be used in a number of problems in electromagnetics, acoustics, and quantum mechanics. The evaluation of the periodic sums with dynamic (Helmholtz potential) PGF can be used in the context of electromanetic and acoustic integral equations for characterizing wave propagation, radiation, scattering, and dispersion diagrams for periodic arrays. The evaluation of the periodic sums with NPSP PGF can be used for modeling infinite periodic arrays or mimic infinite domains in electro-/magneto-statics, micromagnetics, and density functional theory. The evaluation of the periodic sums with static PGF with a phase shift can be used to calculate dispersion diagrams in micromagnetics, e.g., spin waves or linearized density functional theory.